How to find mean differences6) The following SPSS outputs are from a test to compare which means are different from others in a test designed to test mearences from the given table?
Answer #1

As your question is how to find mean difference, so I am explaining you the process..

Mean difference is just the difference of means between two groups..

Foe example:

Group 1 Group 2
1 6
2 7
3 8
4 9
5 10

Mean of group 1 is: (1 + 2 + 3 + 4 + 5)/5 = 3

Mean of group 2 is (6 + 7 + 8 + 9 + 10)/5 = 8

Mean Difference between Group 1 and 2 is: (3 - 8) = -5

and..

Mean Difference between Group 2 and 1 is: (8 - 3) = 5

So for the above output, the mean difference in the first raw is between 1 and the remaining .

Suppose,

For 1 and 2, as shown in the output, the M.D. is 0.333 (First Row)... This means that the mean of Group 1 is more than that of 2 and that is why the difference is positive..

Similarly,

For 2 and 1, as shown in the output, the M.D. is -0.333 (Second Row)... This means that the mean of Group 2 is less than that of 1 and that is why the difference is negative..

It is obvious that if mean of one group is more than the other, then the difference between the highest and the lowest is positive and in reverse case, it is negative...